# Mass Shootings, Community Mobility, and the Relocation of Economic Activity

### v1

@@ 5 Conclusion and Discussion
Our paper discusses an important, yet unexplored, cost of mass shootings: welfare loss due to displacement of economic activities. Visits to nearby POIs decrease after a mass shooting by about 30% with respect to the baseline, while visits to farther-away POIs significantly increase. The impact is not only substantial in size but also long-lasting and pervasive across a variety of location types. Evidence suggests that there is relocation of economic activity, where individuals substitute away from POIs near the shooting site, particularly for places that are easily substitutable.
@@
A key driver of relocation is the psychological effects of these events. The impact of mass shootings is stronger when they occur in safer areas, when they receive more attention from the media, and for POIs that previously received many visitors. Relocation of activities carries great economic implications, as it would involve friction costs and efficiency losses due to distortions in optimal choices of time and location. This could, in turn, negatively impact other economic outcomes, such as earnings and productivity.
For future research, it would be desirable to estimate the monetary costs of mass shootings to local businesses using POI-level transaction data, if available. Also, the impact identified in this paper is a causal effect driven by unpredictable and exogenous events. Recreational spaces, educational facilities, grocery stores, and temporary accommodation places appear to be among the categories most affected by mass shootings. Therefore, future studies may leverage the temporal and spatial variation of activities in these categories to uncover important relationships between other economic variables.
B.2 Nonparametric Estimation using Splines
We estimate the change in weekly visits for each week number as a function of “continuous” distance from the shooting site using a spline function (piecewise polynomial). We use a cubic spline with four knots, where the quantile rank for distance, i.e., the proportion of distances that are smaller than a given distance, serves as the running variable and the four knots are set at 0.25, 0.50, 0.75, and 1. Changes are estimated with respect to the week immediately before the shootings (the baseline week). The x-axis and y-axis of the graphs in this section are “Distance Percentile” and “Change in Weekly Visits,” respectively, and therefore, the graphs plot the change in foot traffic to each distance percentile, compared to the foot traffic in week -1. The gray areas in the plots represent 95% confidence intervals and the standard errors are clustered at the POI level.
The purpose of this exercise is two-fold. First, it shows that the impact of mass shootings on mobility vanishes beyond five miles from the shooting, supporting our choice of a five-mile radius as the cutoff for sample selection. Second, the patterns here align with those reported in Section 4, demonstrating that our results are robust to the cut-offs used to define our distance bands.
Figure B.2: Change In Weekly Visits Over Time and Distance
(a) Including Sprees
(b) Excluding Sprees
Online Appendices C Distribution of POIs
C.1 Relationship Between Raw Distance and Distance Quantiles
Figure C.1: Relationship Between Raw Distance and Distance Quantiles
C.2 Distribution of POIs by POI Category
Figure C.2: Distribution of POIs over Distance, by POI Category
Online Appendices D Impact by POI Type and POI Substitutability
D.1 Impact by POI Type
Figure D.1: Evolution of Weekly Visits By Category
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. “Near”/“Far” refers to POIs below/above median distance from the shooting location.
D.2 Impact by POI Substitutability: 4-quantile distance bands
Figure D.2-1: Change In Weekly Visits by Ease of Substitution (Using Common Categories)
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Figure D.2-2: Change In Weekly Visits by Ease of Substitution (Using Competition Level)
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Online Appendices E Degree of Traumatic Stress and Magnitude of Impact: 4-Quantile Distance Bands
Figure E-1: Evolution of Weekly Visits by Preexisting Crime Rate
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Figure E-2: Evolution of Weekly Visits by Media Coverage
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Figure E-3: Evolution of Weekly Visits by Visitor Density
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.

### v2

@@ 5 Conclusion
Our paper discusses an important, yet unexplored, cost of mass shootings: welfare loss due to the displacement of economic activity. Visits to nearby POIs decrease a few weeks after a mass shooting by about 26% with respect to the baseline, while farther-away POIs experience noticeable increases. The impact is not only substantial in size but also long-lasting and pervasive across a variety of location types. Evidence suggests that there is relocation of economic activity, where individuals substitute away from POIs near the shooting site, particularly for places that are easily substitutable.
@@
A key driver of displacement is the psychological effects of these events. The impact of mass shootings is stronger when they occur in safer areas, when they receive more attention from the media, and for POIs that previously received many visitors. Displacement of activities carries great economic implications, as it would involve friction costs and efficiency losses due to distortions in optimal choices of time and location. This could, in turn, negatively impact other economic outcomes, such as earnings and productivity.